OpenStudy (anonymous):

What is the Area of a regular Hexagon with a side of 6 cm? ( Thousandths )

OpenStudy (anonymous):

A = 1/2aP A = 1/2(a)(36) A = 18a A = 18(3√3) A = 54√3 I'm assuming that you know how to find an apothem and how to plug in √3 into the calculator?